### 02 Mar 2024 v2

# Superconducting Electronic Device Response Linearity Calculation Method Using Discrete Fourier Analysis

Abstract
A method for estimating the linearity of superconducting electronic devices is discussed. The method allows the obtaining of expected spurious-free dynamic range values based on the results of preliminary numerical modeling and solving optimization problems. A description of the method and an analysis of measurement errors are provided.
SFDR, superconducting electronics, numerical simulation, linearity, device modeling
I Introduction
Despite the nonlinear nature of the Josephson effect, it is required for many tasks, such as signal reception, analog-to-digital conversion and signal amplification, to achieve a certain level of linearity. When we talk about the experimental evaluation of the linearity of a physical device, standard methods can be used, such as the single-tone or two-tone methods for determining spurious-free dynamic range (SFDR) [1]. Nevertheless, it is often necessary to have an estimate of system linearity at design stage. Depending on the specific task, authors may use Pearson’s chi-squared test for linear fitting [2] or the first derivative method [3] to evaluate the deviation of the signal under study from an optimal linear signal. These methods are useful for optimization problems, but the results are difficult to compare to subsequent experimental values.
This article consistently considers a method for numerically evaluating the linearity of the output signal. This allows obtaining linearity in terms of SFDR. The method has been successfully applied by our research group in solving optimization problems with differential quantum cells, bi-SQUIDs, electrically small antennas (ESA), and other devices [4, 5, 6].
II Spurious-free dynamic range as a measure of linearity
In the following method, we will use the definition of device linearity, which is defined as SFDR. The linearity of the device’s response can be calculated by applying an input signal in the form of a one-tone signal at frequency $\omega_{1}$ or two tones at equal amplitudes at $\omega_{1}$ and $\omega_{2}$. The output signal spectrum is then measured. The formula for calculating the response linearity is:
$\text{Lin}=20\cdot\log\left(a_{1}/\max{\left\{a_{k}\right\}}\right)$ (1)
where $a$ represents the fundamental tone amplitude and $a_{k}$ is the amplitude of the harmonics of the signal ($k\omega_{1}$ at $k=2,3,\dots$) in a one-tone approach or the intermodulation components at frequencies $n\omega_{1}\pm m\omega_{2}$ ($n$ and $m$ are integers) in the two-tone analysis technique, when $k=n+m=2,3,\dots$.
In fact, it should be noted that SFDR is primarily a measure of signal purity, and not just linearity. This is because it takes into account both noise (unwanted spurious signals) and distortion caused by nonlinearity, and thus essentially represents the effective dynamic range of a system. However, unlike noise, which can have a lower RMS value with a decreasing frequency band, nonlinear distortion places fundamental limitations on the SFDR, making it a measure of linearity in itself.
III Algorithm
The main idea of the method for determining the linearity from the calculated data is quite simple. The input data for the method is the output signal of the device $V_{\text{signal}}(F)$ (where $F$ is a parameter [^1] In real devices, this could be, e.g., an external magnetic flux.) calculated by some program (PSCAN [7, 8], WRSpice [9], etc.), and the result of applying this method should be the linearity of this signal in terms of SFDR.
Let us consider this method in details (see the flowchart in figure 1):
Figure 1: Flowchart for linearity calculation algorithm. $V_{\text{signal}}(F)$ – original discrete signal (calculated by outer program), $\mathfrak{V}_{\text{signal}}(F)$ – continuous signal, $\mathfrak{V}_{one-tone}(t)$ – output signal of device, when a one-tone signal applied as input, $\left\{\left(V_{\text{one-tone}}\right)_{n}\right\}$ – FFT-ready array, $\left\{a_{k}\right\}$ set of amplitudes of harmonics. Dashed lines represents discrete functions (or arrays), solid – continuous functions.
  1. 1.
    Suppose we have a calculated signal $V_{\text{signal}}(F)$, which is a set of $N$ points of the form $\left(V_{\text{signal}}\right)_{n}(F_{n})$. In general, the $F_{N}$ may not be distributed evenly. However, for simplicity, we will assume that they were calculated with a constant step $\Delta F$. Note that this signal does not yet contain the time-dependence. This is the device’s response function and not the quasi-harmonic signal under investigation. The modification of the method in the event that the simulation results in time dependence $V_{\text{signal}}(t)$ will be considered in the Appendix A.
  2. 2.
    Before applying a sinusoidal signal to our response function, we must move from a discrete consideration to a continuous one. Since we postulate that the response is accurately measured at points $F_{n}$ (with an allowance for the relative error $\varepsilon$), it is logical to choose an interpolation method to describe the system between points $F_{n}$ and $F_{n+1}$. Moreover, the use of approximation methods would lead to an obviously incorrect result. Indeed, if we approximate our response with a polynomial $P(x)$ of degree $m$, then further analysis would show the linearity of this polynomial rather than the original response. The only way to determine the linearity of the original response would be through the accuracy of the approximation, which is precisely what we wished to avoid. Therefore, for reconstructing the response, we will employ an interpolation technique using polynomials of degree $m$. This degree $m$ should be sufficient for the intended description of our system. For most SQUID systems, it is sufficient to choose $m=3\text{ to }6$. Denote this reconstructed response as $\mathfrak{V}_{\text{signal}}(F)$.
  3. 3.
    Using $\mathfrak{V}_{\text{signal}}(F)$, we can obtain the device’s output signal when a one-tone signal with a certain frequency $\omega$ is applied to its input: $\mathfrak{V}_{one-tone}(t)=\mathfrak{V}\left(A\sin\left(\omega t\right)+A_{0}\right)$, where $A$ and $A_{0}$ are some constants (amplitude and offset, respectively). The choice of these constants is determined entirely by the choice of the part of the response we wish to investigate. The choice of frequency (in relative units) depends on the number of points $2^{P}$ used for further Fourier analysis. As we have the flexibility to change each of these quantities, we can set the frequency arbitrarily (if we assume that the response does not depend on frequency; otherwise, we need to use the modified method described in the Appendix A). Note that $\mathfrak{V}_{one-tone}(t)$, defined in this way, is definitely continuous in $t$.
  4. 4.
    From the dependence $\mathfrak{V}_{one-tone}(t)$, we can derive $2^{P}$ points $\left\{\left(V_{\text{one-tone}}\right)_{n}\right\}$ in the interval $t\in\left[0;T\right]$. $T$ should be large enough to include many periods, i.e. $T\gg 1/\omega$. The resulting array $\left\{\left(V_{\text{one-tone}}\right)_{n}\right\}$ can then be processed by Fast Fourier Transform (FFT), which will yield a set of amplitudes $\left\{a_{k}\right\}$, where $a_{1}$ represents the amplitude of the fundamental tone and $a_{k}$ ($k\geq 2$) is the amplitude of harmonic components of the signal.
  5. 5.
    Finally, using the array $\left\{a_{k}\right\}$, we can determine the linearity by the formula (1).
IV Method discussion
IV.1 Requirements for the input data of the method (signal)
We assume that the input data ($V_{\text{signal}}$) were obtained from some program for modeling superconducting electronics devices. To make the method work, it is necessary to ensure that the simulated signal adequately describes the response function of the device. The term “adequately” can be understood as follows [10]: the number of points in the simulated signal allows us to make an unambiguous assumption about the amplitude and shape of that signal.
IV.2 FFT accuracy
FFT methods are characterized by high accuracy. Moreover, when using accurate FFT methods (such as, e.g., the Cooley-Tukey FFT algorithm [11]), the introduced error depends only on the floating-point accuracy $\varepsilon_{\text{float}}$ and is approximately equal to $\varepsilon_{\text{float}}\sqrt{P}$. This error can be estimated to be $10^{-10}$. Thus, the introduced error is considered to be insignificant.
IV.3 The issue of selecting FFT parameters
We need to choose three parameters when processing an FFT signal: the frequency $\omega$, the number of points $2^{P}$, and the time $T$. By converting from a discrete function to a continuous one, we can avoid the standard problems associated with FFT analysis and instead use a rectangular window. For proper consideration, we just need to formulate the following requirements for $T$, $2^{p}$, and $\omega$: an arbitrary integer number of cycles should fit within $T$, so $T$ should be equal to $M\times 2\pi/\omega$, where $M$ is an integer; and for harmonic detection purposes, it’s reasonable to require that $2^{P}\gg\left(5\text{ to }10\right)\times\omega$.
IV.4 The issue of the effect of input signal inaccuracy on the obtained linearity value
Taking into account the transfer of inaccuracy (error) of the input signal (the so-called fuzzy data) into the error of the interpolation polynomial is a non-trivial problem [12]. Its solution may be using fuzzy splines [13]. Denote the relative error in the original data by $\varepsilon$ and the resulting error by $\varepsilon_{\text{int}}$.
Another factor affecting the inaccuracy in the obtained linearity value is the transformation of the error $\varepsilon_{\text{int}}$ during the FFT. This resulting error $\varepsilon_{\text{lin}}$ is the result error of our method. Although FFT can be considered as linear, this issue is often solved through numerical experiments, especially for estimating the upper bound [14].
The key idea behind the numerical experiment to estimate $\varepsilon_{\text{int}}$ is as follows: we conduct a large number of measurements of the test signal (a linear signal with random errors with a relative magnitude of $\varepsilon$), obtaining a certain distribution of linearity values. We then consider the relative error $\varepsilon_{\text{lin}}$ as $\sqrt{\sigma^{2}}/M$, where $\sigma^{2}$ is the dispersion and $M$ is the mean value of this distribution.
[…]
Figure 2: Linearity relative error $\varepsilon_{\text{lin}}$ versus relative error $\varepsilon$ of input signal. Data obtained through numerical experiment (the number of measurements is $10^{3}$, the polynomial degree $m=6$ and the number of FFT points is $2^{12}$).
It should be noted that the inaccuracy of the original signal (both numerical noise $\varepsilon$ and specially introduced sources for modeling thermal noise) affect the linearity lowering its value. This is not a method error but rather a reflection by the method of the actual decrease in the values of linearity: the presence of noise affects the amplitude of harmonics and consequently lowers linearity in terms of SFDR.
V Conclusion
[…]
Acknowledgements.
This research was funded by Russian Science Foundation (RSCF) grant no. 19-72-10016-P. N.V.K. is grateful to T. Filippov for his advice to consider this issue in more detail, which led to the writing of this article.
CRediT author statement
Conceptualization: N.V.K. and V.K.K.; Formal analysis: N.V.K.; Funding acquisition: N.V.K.; Investigation: N.V.K.; Methodology: N.V.K. and V.K.K.; Project administration: N.V.K.; Software: N.V.K.; Supervision: N.V.K. and V.K.K.; Validation: N.V.K. and V.K.K.; Writing – original draft: N.V.K.
Conflict of interest statement
The authors declare no conflict of interest. The funders had no role in any part of the manuscript preparation process.
Appendix A Modification of the algorithm if a one-tone signal is present in the initial signal.
Let us consider a modified algorithm in which the initially signal already contains an applied one-tone signal.
  1. 1.
    Suppose we have a calculated signal $V_{\text{one-tone}}(t)$, which is a set of $N$ points of the form $\left(V_{\text{one-tone}}\right)_{n}(t_{n})$. In general, the $t_{N}$ may not be distributed evenly, however, for simplicity, we will assume that they were calculated with a constant step $\Delta t$.
  2. 2.
    Although we already have a time-dependent signal, we still need to convert our signal into a continuous one. Similarly to the original algorithm, we reconstruct the signal between points $t_{n}$ and $t_{n+1}$ using interpolation methods, denote the reconstructed signal as $\mathfrak{V}_{\text{one-tone}}(t)$.
  3. 3.
    Continuation of algorithm is the same:
    1. (a)
      From the dependence $\mathfrak{V}_{one-tone}(t)$, we can derive $2^{P}$ points $\left\{\left(V_{\text{one-tone}}\right)_{n}\right\}$ in the interval $t\in\left[0;T\right]$, where $T$ must be large enough to include many periods, i.e. $T\gg 1/\omega$. The resulting array $\left\{\left(V_{\text{one-tone}}\right)_{n}\right\}$ can then be processed by Fast Fourier Transform (FFT), which will produce a set of amplitudes $\left\{a_{k}\right\}$, where $a_{1}$ represents the amplitude of the fundamental tone and $a_{k}$ ($k\geq 2$) is the amplitude of harmonics of the signal.
    2. (b)
      Finally, using the array $\left\{a_{k}\right\}$, we can determine the linearity using the formula (1).
Since the methods used are the same (only their order has changed), all the reasoning about the accuracy of the method presented in section IV remains unchanged as well, except for the considerations about selecting FFT parameters. A more complex consideration of using windows is beyond the scope of this appendix. However, taking into account the known frequency of the one-tone signal, we can always choose a time range $T$ such that an integer number of periods fit into it.